OpenStudy (anonymous):

solve the exponential equation (round to four decimal places) 3(7-11^5x-2)=10

OpenStudy (anonymous):

\[3(7-11^{5x-2})=10\]

OpenStudy (anonymous):

divide by 3

OpenStudy (anonymous):

\[7 - 11^{5x-2} = \frac{10}{3}\]

OpenStudy (anonymous):

\[11^{5x -2} = 7 - \frac{10}{3} \]

OpenStudy (anonymous):

Oh sorry take Log with Base 11

OpenStudy (anonymous):

x = 2 - log3 ----------- 5

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

is it wrong ?

OpenStudy (anonymous):

\[x = \frac{ 2 - \log_{11} 3}{5}\]

OpenStudy (anonymous):

my answer was 0.8090 which is says isn't right

OpenStudy (anonymous):

.3